# Deep Hedging with Options

### v2

@@ Abstract
We propose an enhanced deep hedging framework for index option portfolios, grounded in a realistic market simulator that captures the joint dynamics of S&P 500 returns and the full implied volatility surface. Our approach integrates surface-informed decisions with multiple hedging instruments and explicitly accounts for transaction costs. The hedging strategy also considers the variance risk premium embedded in the hedging instruments, enabling more informed and adaptive risk management. In this setting, state-dependent no-trade regions emerge naturally, improving rebalancing efficiency and hedging performance. Tested across simulated and historical data from 1996 to 2020, our method consistently outperforms traditional delta and delta-gamma hedging, demonstrating superior adaptability and risk reduction.
@@ 1   Introduction
Hedging decisions are inherently tied to the information available at the time those decisions are made. Traditionally, most approaches rely on historical data of the underlying asset to construct hedging strategies. Some studies have gone a step further by incorporating localized information from the implied volatility surface, such as at-the-money or short-term implied volatilities, see for instance Bates, (2005), Alexander and Nogueira, (2007), François and Stentoft, (2021). In this paper, we address the hedging problem for an index option portfolio using a richer set of information—namely, characteristics of the full implied volatility surface. By exploiting the structure of the entire surface, we aim to capture market expectations and dynamics. Our approach explicitly accounts for transaction costs, allowing for a more realistic assessment of hedging performance in practice.
Our hedging strategy is distinctive in several important respects. First, we focus on minimizing the terminal hedging error, as opposed to continuously tracking the portfolio value over time—a practice that often incurs substantial transaction costs. Second, we expand the set of hedging instruments beyond the underlying asset to include another derivative instrument, enabling more flexible and potentially more cost-effective risk management.111 Methodologies such as those proposed by Coleman et al., (2007) and Kélani and Quittard-Pinon, (2017) address these challenges by managing local risk while incorporating standard options as hedging instruments. The latter approach also accounts for transaction costs. While these methods offer valuable insights, the proper integration of market expectations remains an open question. Third, transaction costs are integrated directly into the hedging decision-making process, rather than being considered retrospectively, ensuring that the resulting strategy is better aligned with actual market conditions and trading constraints.
We build on the framework introduced by François et al., (2024), extending it to address the risk management of index option portfolios through the inclusion of an additional hedging instrument. This extension introduces significant challenges, both computational and conceptual. The high-dimensional nature of the problem is tackled using risk-aware reinforcement learning (RL), which is well-suited for sequential decision-making under uncertainty.222 Deep hedging, introduced by Buehler et al., (2019), leverages deep reinforcement learning (DRL) to dynamically adapt to evolving market conditions, capturing both shifting expectations and historical patterns. While this approach has shown remarkable flexibility and adaptability (e.g., Cao et al., (2020), Carbonneau, (2021), Cao et al., (2023)), the training of the neural network requires a market simulator. François et al., (2024) demonstrate that deep hedging strategies can effectively mitigate transaction costs while incorporating information from the implied volatility (IV) surface. Their study, however, is limited to hedging European options using only the risk-free asset and the underlying asset. The potential benefits of expanding the hedging set to include additional instruments, alongside IV-informed policies, remain unexplored. However, incorporating options as hedging instruments adds complexity. First, options are costly to trade and these costs are managed partly through dynamic selection of positions in the hedging instruments, and partly through the introduction of a no-transaction region that limits unnecessary trading.333 No-trade regions, which mitigate the impact of transaction costs, have been extensively studied in the portfolio optimization literature. Constantinides, (1986) first introduced the idea that proportional transaction costs give rise to such regions—a concept further developed by Davis and Norman, (1990) and Balduzzi and Lynch, (1999), who emphasized portfolio allocation over rebalancing costs. In the hedging context, optimal rebalancing based on delta variations has been explored by Henrotte, (1993), Toft, (1996), and Martellini and Priaulet, (2002). Hodges and Neuberger, (1989) and Clewlow and Hodges, (1997) examine hedging within a utility-maximization framework. The optimal hedging strategy consists of no-trade bands around delta, whose width depends on the hedger’s risk aversion. Second, their inclusion increases the dimensionality of the action space, complicating the learning process. Third, to ensure that the RL agent learns a true hedging strategy rather than engaging in speculative behavior, we introduce penalty terms in the reward function that discourages excessive risk-taking. This design helps steer the agent toward strategies that align with the core objective of minimizing portfolio risk in a realistic trading environment.
As in François et al., (2024), we incorporate the full dynamics of both the implied volatility surface and S&P 500 returns into information available to the RL agent. Our study is distinctive in that this enriched informational context allows the agent to learn and adapt to the time-varying variance risk premium, which drives the cost of hedging. Such information partially explains departure from the classic delta-gamma hedging strategy; the impact of risk premia manifests in the long run and is thus not captured by myopic Greeks-based approaches.
Our numerical results demonstrate that reinforcement learning (RL)-based hedging strategies consistently outperform classical delta and delta-gamma approaches across a range of risk measures. In the absence of transaction costs, RL agents achieve substantial reductions in mean squared error (MSE), with risk reduced by a factor of six compared to delta hedging. The inclusion of an additional hedging instrument, such as an at-the-money call option, further enhances performance across all metrics. When transaction costs are introduced, RL strategies remain effective, dynamically adjusting their rebalancing behavior via learned no-trade regions. They achieve lower risk at similar or reduced cost compared to benchmark methods. Moreover, backtesting on historical market data confirms the robustness of RL agents, highlighting their ability to manage risk and maintain profitability in real-world conditions.
The outperformance of RL approaches can be explained through a couple of reasons. First, RL strategies rely on smaller and more frequent trades. Such more progressive rebalancing actions reduces the number of instances where large trades need to be unwound shortly after. Second, the early-stage divergence between RL and delta-gamma positions can be attributed to the RL agent’s deliberate limitation of short exposure to the volatility risk premium embedded in the option used for hedging.
@@
The paper is organized as follows. Section 2 frames the hedging problem in terms of a deep reinforcement learning framework. Section 3 provides the components of the market simulator. Section 4 presents the numerical results. [^4] The Python code to replicate the numerical experiments from this paper can be found at the following link: https://github.com/cpmendoza/deep-hedging_with_options.git. Section 5 concludes.
@@ 2.1   The hedging problem
The goal is to hedge a short position in a portfolio of contingent claims written on the same underlying asset, $S$, over the hedging period $0,\ldots,T$. The time-$t$ market value of the portfolio is expressed as $\mathcal{P}_{t}=\Psi_{t}(S_{t})$ for some function $\Psi_{t}$. For illustrative purposes, we use a European straddle portfolio with maturity $T$ in our numerical examples. In this case, the value $\mathcal{P}_{T}$ represents the portfolio’s terminal payoff, which is given by the mapping $\Psi_{T}(S_{T})=\max(S_{T}-K,0)+\max(K-S_{T},0)$ with $K$ being the strike price.
The hedging strategy involves managing a self-financing portfolio composed of the risk-free asset, the underlying asset, and a hedging option. Specifically, the hedging option is a European option on the same underlying asset with a longer maturity $T^{*}>T$. The strategy is represented by the predictable process $\{\phi_{t}\}_{t=1}^{T}$, with $\phi_{t}=(\phi_{t}^{(r)},\phi_{t}^{(S)},\phi^{(\text{O})}_{t})$, where $\phi_{t}^{(r)}$ is the cash held at time $t-1$ and carried forward to the next period. Moreover, $\phi_{t}^{(S)}$ and $\phi^{(\text{O})}_{t}$ are respectively the number of shares of the underlying asset $S$ and the number of hedging options in the hedging portfolio, both held during the interval $(t-1,t]$. The time-$t$ hedging portfolio value is
@@
where $\text{O}_{t}(T^{*})$ is the time-$t$ hedging option value, $\Delta=\frac{1}{252}$ represents the time increment in years, $r_{t}$ is the time-$t$ annualized continuously compounded risk-free rate and $q_{t}$ is the annualized underlying asset dividend yield, both on the interval $(t-1,t]$. To account for transaction costs the self-financing condition entails that for $t=0,\ldots,T-1$,
@@
where $\kappa_{1}$ and $\kappa_{2}$ represent the proportional transaction cost rates for the underlying asset and the hedging option, respectively. In a practical financial context, transaction costs for options are typically higher than those for the underlying asset. Consequently, we assume $\kappa_{1}<<\kappa_{2}$.
@@
Each time-$t$ action $\phi_{t+1}$ is a function of currently available information on the market: $\phi_{t+1}=\tilde{\phi}(X_{t})$ for some function $\tilde{\phi}$ where $X_{t}$ is the state variables vector. Due to Equation (1), $\phi_{t+1}^{(r)}$ is fully determined when $\phi_{t+1}^{(S)}$ and $\phi^{(\text{O})}_{t+1}$ are specified, and as such the time-$t$ action to be chosen is $(\phi_{t+1}^{(S)},\phi^{(\text{O})}_{t+1})$.
@@ 2.2   Reinforcement learning and deep hedging
The problem described in Equation (2) is addressed by directly estimating the policy function (the investment strategy $\tilde{\phi}$) using a policy gradient method. This approach leverages a parametric representation of the policy function through an Artificial Neural Network (ANN). Specifically, a parameter vector $\theta$ is introduced to define the policy $\tilde{\phi}$, which is optimized to minimize the risk measure $\rho$ applied to the hedging error at maturity. Representing the policy generated by the ANN as $\tilde{\phi}_{\theta}$, the hedging strategy is defined as $\phi_{t+1}=\tilde{\phi}_{\theta}(X_{t})$. Problem (2) can therefore be approximated as
@@
Given the inherent continuity of ANNs, the mapping $\phi_{t+1}=\tilde{\phi}_{\theta}(X_{t})$ may lead to frequent small adjustments in the hedging position, potentially increasing long-term transaction costs. To mitigate this effect, we introduce a no-trade region, within which there is no rebalancing.
At time $t$, the no-trade region is determined by the distance between the current portfolio position, $\phi_{t}$, and the next position proposed by the ANN, $\tilde{\phi}_{\theta}(X_{t})$. Specifically, rebalancing occurs only if the cumulative deviation in positions across hedging instruments exceeds a threshold $l$:
$\phi_{t+1}$ =$=\begin{cases}\phi_{t},&\text{if }\lvert\phi_{t}^{(S)}-\tilde{\phi}_{\theta}^{(S)}(X_{t})\rvert+\lvert\phi^{(\text{O})}_{t}-\tilde{\phi}_{\theta}^{(\text{O})}(X_{t})\rvert\leq l,\\[8.0pt] \tilde{\phi}_{\theta}(X_{t}),&\text{otherwise}.\end{cases}$ (4)
This formulation expresses the no-trade region in terms of the number of shares of option contracts, providing a measure of the distance at which rebalancing becomes cost-effective, capturing the trade-off between transaction costs and maintaining proximity to the desired portfolio adjustments. Indeed, when rebalancing actions proposed by the neural network are minor, they are not implemented because (i) this only leads to a small misalignment with the ideal hedging positions and (ii) this allows avoiding transaction costs. [^5] We tried other specifications for the no-trade region (for instance explicitly capturing transaction cost amounts), with results being qualitatively similar. In this framework, both the ANN parameters $\theta$ and the rebalancing threshold $l$ are treated as learnable parameters, allowing the model to jointly optimize the size of rebalancing actions and decisions of whether or not to rebalance.
As shown in François et al., (2024), the policy $\tilde{\phi}_{\theta}$ may inadvertently incorporate speculative elements, such as doubling strategies, where agents continuously increase their exposure in an attempt to recover successive losses. Such strategies are undesirable as they deviate from sound risk management principles. To prevent this problem, we introduce a soft tracking error constraint that penalizes the network during training if the time-$t$ tracking error,
$\xi_{t}^{(\tilde{\phi}_{\theta},\,l)}=\mathcal{P}_{t}-V_{t}^{(\tilde{\phi}_{\theta},\,l)},$ (5)
exceeds the initial hedging portfolio value at any time $t$. This constraint is defined as:
$SC(\theta,\,l)=\mathbb{P}\left(\max_{t\in\{0,\dots,T\}}\left[\xi_{t}^{(\tilde{\phi}_{\theta},\,l)}\right]>V_{0}\right).$ (6)
This design does not penalize gains, consistent with the asymmetric nature of rational agents. As a result, instead of solving Problem (3), the objective function employed in our approach is
$\mathcal{O}_{\lambda}(\theta,\,l)=\rho\left(\xi_{T}^{(\tilde{\phi}_{\theta},\,l)}\right)+\lambda\,SC(\theta,\,l),$ (7)
where $\lambda$ is a penalty parameter that controls the weight of the soft constraint in the optimization process. Its optimal value is determined independently using a validation set as part of the model selection procedure.
@@ 2.3   Neural network optimization
The RNN-FNN network $\tilde{\phi}_{\theta}(\cdot)$, along with the rebalancing threshold $l$, are optimized with the Mini-batch Stochastic Gradient Descent method (MSGD). This training procedure relies on updating iteratively all the trainable parameters of the optimization problem based on the recursive equations
$\theta_{j+1}=\theta_{j}-\eta_{j}^{(1)}\frac{\partial}{\partial\theta}\hat{\mathcal{O}}_{\lambda}(\theta,\,l),$ (8)
$l_{j+1}=l_{j}-\eta_{j}^{(2)}\frac{\partial}{\partial l}\hat{\mathcal{O}}_{\lambda}(\theta,\,l),$ (9)
where $\eta_{j}^{(1)}$ and $\eta_{j}^{(2)}$ are the learning rates that determine the magnitude of change of parameters per time step. These rates are dynamically adjusted using the Adam optimization algorithm. [^6] Adam is an adaptive learning rate method designed to accelerate training in deep neural networks and promote rapid convergence, as detailed in Kingma and Ba, (2015). Additionally, $\hat{\mathcal{O}}(\theta,\,l)$ is the Monte-Carlo estimate of the objective function defined by Equation (7). Further details can be found in Appendix A.1.
@@ 3   Market simulator
We leverage the JIVR model from François et al., (2023), which models the temporal dynamics of S&P 500 returns and various factors driving the IV surface, along with their interdependencies. The JIVR has the advantage of being data-driven, allowing to replicate multiple realistic shapes of the IV surface encountered in practice. It has been calibrated on an extensive data sample including multiple crises; it can therefore reflect a broad array of economic conditions. Finally, the multi-factor nature of the model leads to a flexible relationship between the underlying asset price and volatility surfaces. Such feature allows reflecting self-contained properties of the option market, consistently with the "instrumental approach" of option pricing detailed in Rebonato, (2005). This section describes the JIVR model.

### v3

@@ Abstract
We propose a deep hedging framework for index option portfolios, grounded in a realistic market simulator that captures the joint dynamics of S&P 500 returns and the full implied volatility surface. Our approach integrates surface-informed decisions with multiple hedging instruments and explicitly accounts for transaction costs. The hedging strategy also considers the variance risk premium embedded in the hedging instruments, enabling more informed and adaptive risk management. Tested on a historical out-of-sample set of straddles from 2020 to 2023, our method consistently outperforms traditional delta-gamma hedging strategies across a range of market conditions.
@@ 1  Introduction
Hedging decisions are inherently tied to the information available at the time they are made. Traditional approaches typically rely on dynamics of the underlying asset, which are estimated on historical data. Some studies extend this framework by incorporating localized information from the implied volatility (IV) surface, such as at-the-money or short-term implied volatilities (Bates,, 2005; Alexander and Nogueira,, 2007; François and Stentoft,, 2021). In this paper, we address the hedging problem for an index option portfolio using a richer set of information—namely, characteristics of the full implied volatility surface. By exploiting the structure of the entire surface, we aim to better capture market expectations and variance dynamics.
Capturing such information at each decision point increases the dimensionality of the state vector. This makes reinforcement learning (RL) a natural choice for identifying optimal hedging strategies. Deep hedging, introduced by Buehler et al., (2019), leverages deep reinforcement learning to dynamically adapt to evolving market conditions, capturing both shifting expectations and historical patterns. While this approach has shown remarkable flexibility and adaptability (e.g., Du et al., (2020), Cao et al., (2020), Carbonneau, (2021), Wu and Jaimungal, (2023), Cao et al., (2023)), the training of the neural network requires a market simulator. François et al., (2024) demonstrate that deep hedging strategies can effectively mitigate transaction costs while incorporating information from the implied volatility (IV) surface. Their study, however, focuses on the hedging of European options using only the underlying asset. The potential benefits of expanding the hedging set to include additional instruments, alongside IV-informed policies, remain unexplored.
We build on the framework introduced by François et al., (2024), extending it to address the risk management of index option portfolios through the inclusion of an additional hedging instrument. This extension introduces significant challenges, both computational and conceptual. First, the state vector requires further information about the additional hedging instrument and the portfolio to be hedged. Second, to ensure that the RL agent learns a true hedging strategy rather than engaging in speculative behavior, we introduce penalty terms in the reward function that discourages excessive risk-taking. This design helps steer the agent toward strategies that align with the core objective of minimizing portfolio risk in a realistic trading environment.
Our study is distinctive in that it simultaneously leverages rich information derived from the implied volatility surface and its dynamics, explicitly accounts for transaction costs—which are particularly significant when trading options—and departs from traditional portfolio tracking by adopting a global hedging objective focused on minimizing terminal hedging error.
Numerical results for hedging a short position on a straddle show that all our RL algorithms consistently and substantially outperform the practitioner’s delta and delta-gamma approaches. In some cases, the RL agent relying only on the underlying as the hedging instrument even outperforms delta-gamma hedging: this happens in particular in the presence of transaction costs when a tail risk performance metric is considered.
The outperformance of RL approaches can be attributed to several factors. First, RL strategies typically rely on smaller trades. This more gradual rebalancing reduces the likelihood of having to unwind large positions shortly after they are established. Second, the early-stage divergence between RL and delta-gamma positions reflects the RL agent’s deliberate efforts to limit short exposure to the variance risk premium embedded in the option used for hedging. Thanks to our enriched informational state vector, the RL agent learns and adapts to the time-varying variance risk premium, which is a key driver of hedging costs. The impact of risk premia materializes over the long term and is therefore not captured by myopic Greeks-based approaches.
As the model is trained on market data from 1996 to 2020, we use recent option data from 2021 to 2023 to evaluate whether our trained RL algorithm maintains its performance out-of-sample. For this backtesting study, we introduce a new benchmark: the RL algorithm without IV information. We demonstrate the superiority of the RL algorithms with the full information over both the practitioners’ delta-gamma strategy and the RL algorithms with limited information. The RL algorithms without IV information do not outperform the practitioners’ delta-gamma approach in terms of mean squared hedging errors. These results highlight the importance of feeding relevant market information to the RL hedging agent.
@@
The paper is organized as follows. Section˜2 frames the hedging problem in terms of a deep reinforcement learning framework. Section˜3 provides the components of the market simulator. Section˜4 presents the numerical results. [^1] The Python code to replicate the numerical experiments from this paper can be found at the following link: https://github.com/cpmendoza/deep-hedging_with_options.git. Section˜5 presents the out-of-sample backtesting results. Section˜6 concludes.
@@ 2.1  The hedging problem
The goal is to hedge a short position in a portfolio of contingent claims written on the same underlying asset, $S$, over the hedging period $0,\ldots,T$. The time-$t$ market value of the portfolio is denoted $\mathcal{P}_{t}$. For illustrative purposes, our numerical examples use a European straddle portfolio with maturity $T$. In this case, the value $\mathcal{P}_{T}$ represents the portfolio’s terminal payoff, which is given by the mapping $\Psi_{T}(S_{T})=\max(S_{T}-K,0)+\max(K-S_{T},0)$ with $K$ being the strike price.
The hedging strategy involves managing a self-financing portfolio composed of the risk-free asset, the underlying asset, and a hedging option. Specifically, the hedging option is a European option on the same underlying asset with a longer maturity $T^{*}>T$. The strategy is represented by the predictable process $\{\phi_{t}\}_{t=1}^{T}$, with $\phi_{t}=(\phi_{t}^{(r)},\phi_{t}^{(S)},\phi^{(\text{O})}_{t})$, where $\phi_{t}^{(r)}$ is the cash held at time t1t-1 and carried forward to the next period. Moreover, $\phi_{t}^{(S)}$ and $\phi^{(\text{O})}_{t}$ are respectively the number of shares of the underlying asset $S$ and the number of hedging options in the hedging portfolio, both held during the interval (t1,t](t-1,t]. The time-$t$ hedging portfolio value is
@@
where $\text{O}_{t}(T^{*})$ is the time-$t$ hedging option value, $\Delta=\frac{1}{252}$ represents the time increment in years, $r_{t}$ is the time-$t$ annualized continuously compounded risk-free rate and $q_{t}$ is the annualized underlying asset dividend yield, both on the interval (t1,t](t-1,t]. To account for transaction costs the self-financing condition entails that for $t=0,\ldots,T-1$,
@@
where $\kappa_{1}$ and $\kappa_{2}$ represent the proportional transaction cost rates for the underlying asset and the hedging option, respectively. Transaction costs for options are typically higher than those for the underlying asset. Consequently, we assume $\kappa_{1}<<\kappa_{2}$.
@@
Each time-$t$ action $\phi_{t+1}$ is a function of currently available information on the market: $\phi_{t+1}=\tilde{\phi}(X_{t})$ for some function $\tilde{\phi}$ of the state variables vector $X_{t}$. Due to Equation (1), $\phi_{t+1}^{(r)}$ is fully determined when $\phi_{t+1}^{(S)}$ and $\phi^{(\text{O})}_{t+1}$ are specified, and as such the time-$t$ action to be chosen is $(\phi_{t+1}^{(S)},\phi^{(\text{O})}_{t+1})$.
@@ 2.2  Reinforcement learning and deep hedging
The problem described in Equation (2) is addressed by directly estimating the policy function (the investment strategy $\tilde{\phi}$) using a policy gradient method. This approach leverages a parametric representation of the policy function through an Artificial Neural Network (ANN). Specifically, the policy $\tilde{\phi}$, governed by a parameter vector $\theta$, is optimized to minimize the risk measure $\rho$ evaluated at the terminal hedging error. Representing the policy generated by the ANN as $\tilde{\phi}_{\theta}$, the hedging strategy is defined as $\phi_{t+1}=\tilde{\phi}_{\theta}(X_{t})$. Problem (2) can therefore be approximated as
@@
Given the inherent continuity of ANNs, the mapping $\phi_{t+1}=\tilde{\phi}_{\theta}(X_{t})$ may lead to frequent small adjustments in the hedging position, potentially increasing long-term transaction costs. To mitigate this effect, we introduce a no-trade region, within which there is no rebalancing. In practice, the no-trade region has a negligible impact on the performance of the ANN, but it improves the results of our benchmark strategies. Further details are provided in Appendix˜A.
As shown in François et al., (2024), the policy $\tilde{\phi}_{\theta}$ may inadvertently incorporate speculative elements, such as doubling strategies, where agents continuously increase their exposure in an attempt to recover successive losses. Such strategies are undesirable as they deviate from sound risk management principles. To prevent this problem, we introduce a soft tracking error constraint
$SC(\theta)=\mathbb{P}\left(\max_{t\in\{0,\dots,T\}}\left\{\xi_{t}^{\tilde{\phi}_{\theta}}\right\}>V_{0}\right)$ (4)
that penalizes the network during training if the time-$t$ tracking error,
$\xi_{t}^{\tilde{\phi}_{\theta}}=\mathcal{P}_{t}-V_{t}^{\tilde{\phi}_{\theta}},$ (5)
exceeds the initial hedging portfolio value at any time $t$. This design does not penalize gains, consistent with the asymmetric nature of rational agents. As a result, instead of solving Problem (3), the objective function employed in our approach is
$\mathcal{O}(\theta;\,\lambda)=\rho\left(\xi_{T}^{\tilde{\phi}_{\theta}}\right)+\lambda\,SC(\theta),$ (6)
where $\lambda$ is a hyperparameter that controls the soft constraint weight in the optimization process. It is determined independently using a validation set during the model selection procedure.
@@ 2.3  Neural network optimization
The RNN-FNN network $\tilde{\phi}_{\theta}(\cdot)$ is optimized with the Mini-batch Stochastic Gradient Descent method (MSGD). This training procedure relies on updating iteratively all the trainable parameters of the optimization problem based on the recursive equations
$\theta_{j+1}=\theta_{j}-\eta_{j}\frac{\partial}{\partial\theta}\hat{\mathcal{O}}(\theta;\,\lambda),$ (7)
where $\eta_{j}$ are the learning rates that determine the magnitude of change of parameters per time step. These rates are dynamically adjusted using the Adam optimization algorithm.222 Adam is an adaptive learning rate method designed to accelerate training in deep neural networks and promote rapid convergence, as detailed in Kingma and Ba, (2015). Additionally, $\hat{\mathcal{O}}(\theta;\,\lambda)$ is the Monte-Carlo estimate of the objective function defined by Equation (6). Further details can be found in Appendix˜B.
@@ 3  Market simulator
We leverage the JIVR model from François et al., (2023), which captures the temporal dynamics of S&P 500 returns alongside the key drivers of the IV surface, while accounting for their interdependencies. The JIVR framework works with interpretable factors and enables the replication of a wide range of realistic IV surface shapes observed in practice. [^3] Other approaches could be pursued to generate IV surface scenarios, such as generative AI models detailed in Chen et al., (2023), Choudhary et al., (2024) and Vuletić and Cont, (2024). The market simulator has been estimated using a daily dataset of observed implied volatilities—covering a broad range of moneyness and time-to-maturity—alongside S&P 500 returns from 1996 to 2020; it can therefore reflect a broad array of market conditions. It captures the self-contained properties of the option market, consistently with the "instrumental approach" of option pricing detailed in Rebonato, (2005).